OpenStudy (anonymous):

parking in a student lot costs $1 for the first half hour and $1.25 for each hour there after. a partial hour is charged the same as a full hour what is the longest time that a student can park in this lot for $10?

OpenStudy (eyust707):

Total cost = 1 + 1.25*(hours after the first half hour)

OpenStudy (eyust707):

we want total cost to equal 10

OpenStudy (eyust707):

10 = 1 + 1.25 *(h)

OpenStudy (eyust707):

9=1.25*h

OpenStudy (eyust707):

9/1.25 = h

OpenStudy (eyust707):

7 hours would cost $9.13

OpenStudy (eyust707):

7.1 hours you would be charged for 8 and youd go over

OpenStudy (eyust707):

7 hours is the longest you can stay if you only have $10

OpenStudy (anonymous):

thank you!!!